PrettyEyes:

Mindy's dog, Rex, weighed 80 pounds a year ago. Now he weighs 92 pounds. What is the percent increase in his weight? A. 15% B. 12% C. 16% D. 13%

Angle:

Let's ask this first: how much did the weight increase?

IsabelFriedmannbells:

its not a or c which one do you think it is?

PrettyEyes:

12

Angle:

ok now we want to use the formula: 100 * (increase) / (original) = (percent increase)

Angle:

so what is 100 * 12 / 80 = ?

PrettyEyes:

100*12 is 1200 and 100*80 is 8000

Angle:

nonono that's a division sign after the 12

PrettyEyes:

i dont get it anymore

Angle:

\[100 \times 12 \div 80 = \]

PrettyEyes:

ohhh ok

PrettyEyes:

its 15

Angle:

you got it! :)

Angle:

that's your answer

PrettyEyes:

ohhhhh thank you

IsabelFriedmannbells:

gj